OpenStudy (anonymous):

Suppose you have a triangle. The opposite side(a.k.a b) = 5 and the degrees = 36. Evaluate the hypotenuse (a.k.a. c)

Directrix (directrix):

sin(36) = 5/c .5877 = 5/c .5877c = 5 c = 8.5065